Sales were 4,000 at a price of $200 but fell to 3,800 when the price was increased to $220. Calculate the percentage change in sales, the percentage change in price, and the price elasticity of demand. (THREE PART)

PLEASE SHOW ALL FORMULAS IN DETAIL
Answer #1

It is given that sales were 4,000 when price was $200. So, the quantity demanded equals Sales/Price = 4,000/200 = 200 units.

Sales fell to 3,800 when the price increased to $220. So, the new quantity demanded equals 3,800/220 = 17.27 units.

Percentage change in price can be calculated by the following formula:

New Price Old Price Old Price x 100

220 200 200 x 100

20 x 100 10 200

So, there is a 10% increase in price.

Percentage change in sales can be calculated by the following formula:

New Sales Old Sales Old Sales x 100

3,800 4,000 x 100 4,000

-200 x 100 5 4,000

So, there is a 5% fall in sales.

The price elasticity of demand can be calculated by the following formula:

New Quantity Demanded Old Quantity Demanded Old Price Old Quantity Demanded New Price Old Price

17.27 20 200 X 220 200 20

2.27 200 X 20 20

35

Therefore, the price elasticity of demand is -1.135.